# RSVG-ZeroOV: Exploring a Training-Free Framework for Zero-Shot Open-Vocabulary Visual Grounding in Remote Sensing Images

### v1

@@ Introduction
Deep learning has been widely applied in remote sensing tasks, significantly improving the performance of applications such as image classification, object detection, and semantic segmentation (Li et al. 2023; Chen et al. 2024a; Shan et al. 2025; Pan et al. 2025). However, these advances are heavily dependent on large-scale annotated data to learn how to recognize and localize objects. Given the broad geographic coverage and the diverse types of objects present in remote sensing imagery, existing datasets cannot fully capture the diversity of real-world objects. As a result, when encountering out-of-distribution objects, existing models may struggle to recognize and localize these objects, limiting their reliability in practical remote sensing scenarios.
@@
Recently, generic foundation models have demonstrated remarkable open-world perception capabilities (Liu et al. 2023; Kirillov et al. 2023; Liu et al. 2024d). Benefiting from pre-training on extensive image-text pairs, these models are able to effectively understand the complex semantic relationships between textual descriptions and visual elements, thereby achieving impressive vision-language alignment.
Some researchers (Yao et al. 2025; Huang, He, and Wen 2025) have begun exploring the potential of foundation models in open-vocabulary category-driven tasks, such as remote sensing object detection and segmentation, relying on predefined object categories (e.g., “road”, “farmland”) to calculate the similarity with image regions, as illustrated in Fig. 1 (left and middle). Furthermore, in many real-world scenarios, users are more interested in locating specific targets. For instance, in urban analysis, planners may seek to identify “the tallest building by the river” or “the factory next to the playground.” Moreover, some objects cannot be represented by a simple category name. For example, temporary roadside parking zones may visually resemble roads but functionally serve as informal parking lots. Such expressions require models to reason about visual attributes, spatial relationships, and functional roles. Conventional open-vocabulary detection or segmentation methods are insufficient for handling such intention-driven tasks. In contrast, zero-shot open-vocabulary remote sensing visual grounding (RSVG) offers a more practical and flexible solution.
@@
Generic Foundation Model. The emergence of large-scale pretraining has fueled the rapid advancement of multimodal foundation models. Among them, visual-language models (VLMs) are designed for cross-modal understanding by aligning visual and textual representations. Early attempts like CLIP (Radford et al. 2021) demonstrated impressive zero-shot generalization by contrastively pretraining on massive image-text pairs. LLaVA (Liu et al. 2023) builds upon CLIP by adding a language decoder Vicuna (Chiang et al. 2023) to support multimodal instruction-following. More recent approaches, such as CogVLM (Wang et al. 2024b), InternVL (Chen et al. 2024b), and Qwen-VL (Wang et al. 2024a), further enhance this paradigm and demonstrate outstanding performance in open-world visual understanding. However, due to their emphasis on high-level semantic reasoning and lack of pixel-level supervision, these models often struggle with fine-grained spatial perception. In contrast, another branch represented by diffusion models (DMs) (Rombach et al. 2022), focuses on visual content generation from text prompts. Recent studies (Pnvr et al. 2023; Xu et al. 2023) reveal that DMs’ hierarchical attention encodes strong structural priors, making their attention maps applicable to downstream tasks like classification, detection, and segmentation. In this study, we conduct systematic exploratory experiments to investigate the potential of integrating VLM and DM models for open-vocabulary RSVG, which will be elaborated in the next section.
@@
Remote Sensing Visual Grounding. RSVG aims to localize objects in remote sensing imagery based on natural language descriptions. Compared to natural scene images, remote sensing data is acquired via satellites and often encounters unique challenges such as scale variation and cluttered backgrounds. Remote sensing referring expression comprehension (RSREC) and segmentation (RSRES) are two typical RSVG tasks. RSREC aims to predict a bounding box corresponding to the given referring expression. The pioneering work (Sun et al. 2022) explores RSREC through a one-stage method based on dense anchors. Recently, transformer-based methods (Zhan, Xiong, and Yuan 2023; Li et al. 2024; Lan et al. 2024; Ding et al. 2024) can effectively capture cross-modal context and achieve better performance, benefiting from the attention mechanism. For RSRES, the goal is to predict a precise pixel-wise binary mask specified by the given referring expression. Recent works (Yuan et al. 2023; Liu et al. 2024c; Dong et al. 2024) improve segmentation by enhancing multi-scale visual-text feature interactions.
Open-Vocabulary Visual Perception. With the rapid development of multimodal foundation models, open-vocabulary learning has achieved remarkable success in natural image domains. Inspired by this progress, recent studies have extended these paradigms to the remote sensing field. For instance, RemoteCLIP (Liu et al. 2024a) improves transferability in downstream tasks by pretraining on extensive remote sensing image-expression pairs. Similarly, VisGT (Pan et al. 2025) and GSNet (Ye, Zhuge, and Zhang 2025) construct large-scale open-vocabulary detection and segmentation datasets, respectively, enabling models to recognize a broader range of object categories. However, curating high-quality annotated datasets for remote sensing remains costly and labor-intensive. To alleviate reliance on manual annotations, several efforts (Zheng et al. 2025) have explored training-free approaches. For example, SegEarth-OV (Li et al. 2025) adapts existing generic models by modifying or removing task-specific components, allowing seamless application to high-resolution remote sensing segmentation. Nevertheless, these methods still rely on predefined category labels during inference and lack the flexibility to understand user-intended objects described in free-form language. As discussed in the introduction, users in practical applications often prefer to locate objects by describing visual attributes, spatial relationships, or functional roles using natural language expressions. To address this limitation, we introduce the task of zero-shot open-vocabulary RSVG, which requires models to flexibly interpret free-form expressions and localize the referred objects. To this end, we propose a training-free framework that integrates attention maps from frozen VLM and DM to facilitate this task.
@@
Guideline 1: generic VLMs exhibit strong generalization ability on unseen remote sensing data. To demonstrate this, we conduct zero-shot experiments on RRSIS-D test set. Specifically, we compare three 7B-scale models, including LLaVA-1.5 (Liu et al. 2024b), Qwen2.5-VL (Bai et al. 2025), and GeoChat (Kuckreja et al. 2024). For RSREC task, we generate bounding box predictions of referred objects using a unified prompt format: [Locate the object referred to by ‘{referring expression}’ and return its box coordinates (x1, y1, x2, y2)]. Notably, GeoChat has been finetuned on 318k image-instruction pairs collected from remote sensing datasets [^2] A portion of GeoChat’s training data shares source images (from DIOR (Li et al. 2020) dataset) with RRSIS-D test set, making the evaluation not strictly zero-shot for GeoChat., enabling it to learn more domain-specific priors. However, as shown in Tab. 1, GeoChat achieves 23.93% Pr@0.5 on RSREC task, noticeably lower than Qwen2.5-VL’s 28.66%. This result suggests that VLMs pretrained on large-scale generic datasets can exhibit strong cross-domain generalization even without access to remote sensing data. For RSRES task, we utilize outputs of VLMs as box prompts for SAM to get pixel-level masks. As we can see, Qwen2.5-VL still outperforms GeoChat by 0.8% in mIoU. These findings highlight the potential of generic VLMs as a solid foundation for zero-shot open-vocabulary RSVG, even without any domain-specific training.
@@
Refine Stage (Optional). Following prior work (Liu et al. 2024c; Dong et al. 2024), we additionally introduce an optional refinement stage to further improve the segmentation mask $\mathbf{M}$. We evaluate the effectiveness of several refinement methods, including DenseCRF and various prompt types for SAM. Our results show that box prompts yield the best performance, whereas point prompts are less effective. Algorithm table, detailed experimental results, and corresponding analyses are provided in Appendix A.
@@
Datasets. We evaluate our method on two benchmarks tailored for RSVG: RRSIS-D (Liu et al. 2024c) and RISBench (Dong et al. 2024). Each sample comprises an RGB image, a referring expression, and a segmentation mask. Both datasets are extended from the RSREC dataset, inheriting its bounding box annotations, which allows for comprehensive evaluation of both REREC and RSRES tasks. Further descriptions can be found in Appendix B.1.
Implementation Details. Our RSVG-ZeroOV offers a zero-shot solution, requiring only an inference process, without the need for any training images or annotations. We employ the pre-trained Qwen2.5-VL (Bai et al. 2025) and Stable Diffusion V1.4 (Rombach et al. 2022) as our VLM and DM, respectively. For Stable Diffusion, we use a guidance scale of 7.5 with 1,000 total diffusion steps, and perform DDIM-based sampling with 20 steps. The DDIM noise schedule follows a scaled linear progression from 0.00085 to 0.012. Following prior works (Li et al. 2024; Liu et al. 2024c), we report mIoU, oIoU and Precision@X with $X\in\{0.3,0.5,0.7\}$. We set KK=7 for seed selection, $\tau$=0.3 for response thresholding, and $\alpha$=0.4 for binarization. Detailed hyperparameter analyses are provided in Appendix C.4. All experiments are conducted on a single RTX4090 GPU.
@@
Comparison Results on RISBench. We further evaluate RSVG-ZeroOV on the more challenging RISBench, which features longer and more semantically complex referring expressions compared to RRSIS-D. As shown in Tab. 6, RSVG-ZeroOV (w/ Refine) demonstrates exceptional performance, establishing new SOTA benchmarks for Pr@0.5 and mIoU at 38.90%/38.87% and 31.03%/31.84% on both tasks, respectively. Notably, the proposed method outperforms the two zero-shot methods, OV-VG and GroundVLP by +18.8%/18.13% and +17.15%/16.27% in oIoU, respectively, indicating its superior ability to capture object structures and preserve overall shape integrity. More visualizations and analyses are provided in Appendix C.2.
@@
Effect of self-attention resolutions. As shown in Tab. 8, we investigate the impact of varying resolutions of self-attention maps. We find that using multiple resolutions [32, 64] leads to the best performance, suggesting that multi-scale self-attention maps not only preserve high-resolution details but also effectively capture contextual semantics. Moreover, more ablation experiments and parameter analyses are provided in Appendix C.4.
@@ Conclusion
In this paper, we introduce RSVG-ZeroOV, a zero-shot framework that integrates a VLM and a DM via attention maps to tackle the open-vocabulary remote sensing visual grounding. Without additional training, we design a novel overview-focus-evolve strategy to progressively enhance attention toward referred objects. Extensive experiments on existing benchmarks demonstrate that RSVG-ZeroOV achieves SOTA performance, outperforming existing weakly-supervised and zero-shot methods. To the best of our knowledge, RSVG-ZeroOV is the first zero-shot framework for remote sensing visual grounding. We hope this framework can potentially inspire researchers to address open-vocabulary remote sensing tasks from a fresh perspective. A discussion of limitations and potential future directions is provided Appendix D.

### v2

@@ Introduction
Deep learning has been widely applied in remote sensing tasks, significantly improving the performance of applications such as image classification, object detection, and semantic segmentation (chen2024rsprompter; li2024show; shan2025ros; pan2025locate). However, these advances are heavily dependent on large-scale annotated data to learn how to recognize and localize objects. Given the broad geographic coverage and the diverse types of objects present in remote sensing imagery, existing datasets cannot fully capture the diversity of real-world objects. As a result, when encountering out-of-distribution objects, existing models may struggle to recognize and localize these objects, limiting their reliability in practical remote sensing scenarios.
@@
Recently, generic foundation models have demonstrated remarkable open-world perception capabilities (liu2023visual; kirillov2023segment; liu2024grounding). Benefiting from pre-training on extensive image-text pairs, these models are able to effectively understand the complex semantic relationships between textual descriptions and visual elements, thereby achieving impressive vision-language alignment.
Some researchers (yao2025remotesam; huang2025zori) have begun exploring the potential of foundation models in open-vocabulary category-driven tasks, such as remote sensing object detection and segmentation, relying on predefined object categories (e.g., “road”, “farmland”) to calculate the similarity with image regions, as illustrated in Fig. 1 (left and middle). Furthermore, in many real-world scenarios, users are more interested in locating specific targets. For instance, in urban analysis, planners may seek to identify “the tallest building by the river” or “the factory next to the playground.” Moreover, some objects cannot be represented by a simple category name. For example, temporary roadside parking zones may visually resemble roads but functionally serve as informal parking lots. Such expressions require models to reason about visual attributes, spatial relationships, and functional roles. Conventional open-vocabulary detection or segmentation methods are insufficient for handling such intention-driven tasks. In contrast, zero-shot open-vocabulary remote sensing visual grounding (RSVG) offers a more practical and flexible solution.
@@
Generic Foundation Model. The emergence of large-scale pretraining has fueled the rapid advancement of multimodal foundation models. Among them, visual-language models (VLMs) are designed for cross-modal understanding by aligning visual and textual representations. Early attempts like CLIP (radford2021learning) demonstrated impressive zero-shot generalization by contrastively pretraining on massive image-text pairs. LLaVA (liu2023visual) builds upon CLIP by adding a language decoder Vicuna (chiang2023vicuna) to support multimodal instruction-following. More recent approaches, such as CogVLM (wang2024cogvlm), InternVL (chen2024internvl), and Qwen-VL (wang2024qwen2), further enhance this paradigm and demonstrate outstanding performance in open-world visual understanding. However, due to their emphasis on high-level semantic reasoning and lack of pixel-level supervision, these models often struggle with fine-grained spatial perception. In contrast, another branch represented by diffusion models (DMs) (rombach2022high), focuses on visual content generation from text prompts. Recent studies (pnvr2023ld; xu2023open) reveal that DMs’ hierarchical attention encodes strong structural priors, making their attention maps applicable to downstream tasks like classification, detection, and segmentation. In this study, we conduct systematic exploratory experiments to investigate the potential of integrating VLM and DM models for open-vocabulary RSVG, which will be elaborated in the next section.
@@
Remote Sensing Visual Grounding. RSVG aims to localize objects in remote sensing imagery based on natural language descriptions. Compared to natural scene images, remote sensing data is acquired via satellites and often encounters unique challenges such as scale variation and cluttered backgrounds. Remote sensing referring expression comprehension (RSREC) and segmentation (RSRES) are two typical RSVG tasks. RSREC aims to predict a bounding box corresponding to the given referring expression. The pioneering work (sun2022visual) explores RSREC through a one-stage method based on dense (ding2024visual). Recently, transformer-based methods (zhan2023rsvg; li2024language; lan2024language; ding2024visual) can effectively capture cross-modal context and achieve better performance, benefiting from the attention mechanism. For RSRES, the goal is to predict a precise pixel-wise binary mask specified by the given referring expression. Recent works (yuan2023rrsis; liu2024rotated; dong2024cross; li2025segearth_r1) improve segmentation by enhancing multi-scale visual-text feature interactions.
Open-Vocabulary Visual Perception. With the rapid development of multimodal foundation models, open-vocabulary learning has achieved remarkable success in natural image domains. Inspired by this progress, recent studies have extended these paradigms to the remote sensing field. For instance, RemoteCLIP (liu2024remoteclip) improves transferability in downstream tasks by pretraining on extensive remote sensing image-expression pairs. Similarly, VisGT (pan2025locate) and GSNet (ye2025towards) construct large-scale open-vocabulary detection and segmentation datasets, respectively, enabling models to recognize a broader range of object categories. However, curating high-quality annotated datasets for remote sensing remains costly and labor-intensive. To alleviate reliance on manual annotations, several efforts (zheng2025instructsam) have explored training-free approaches. For example, SegEarth-OV (li2025segearth_ov) adapts existing generic models by modifying or removing task-specific components, allowing seamless application to high-resolution remote sensing segmentation. Nevertheless, these methods still rely on predefined category labels during inference and lack the flexibility to understand user-intended objects described in free-form language. As discussed in the introduction, users in practical applications often prefer to locate objects by describing visual attributes, spatial relationships, or functional roles using natural language expressions. To address this limitation, we introduce the task of zero-shot open-vocabulary RSVG, which requires models to flexibly interpret free-form expressions and localize the referred objects. To this end, we propose a training-free framework that integrates attention maps from frozen VLM and DM to facilitate this task.
@@
Guideline 1: generic VLMs exhibit strong generalization ability on unseen remote sensing data. To demonstrate this, we conduct zero-shot experiments on RRSIS-D test set. Specifically, we compare three 7B-scale models, including LLaVA-1.5 (liu2024improved), Qwen2.5-VL (Qwen2.5-VL), and GeoChat (kuckreja2024geochat). For RSREC task, we generate bounding box predictions of referred objects using a unified prompt format: [Locate the object referred to by ‘{referring expression}’ and return its box coordinates (x1, y1, x2, y2)]. Notably, GeoChat has been finetuned on 318k image-instruction pairs collected from remote sensing datasets [^2] A portion of GeoChat’s training data shares source images (from DIOR (li2020object) dataset) with RRSIS-D test set, making the evaluation not strictly zero-shot for GeoChat., enabling it to learn more domain-specific priors. However, as shown in Tab. 1, GeoChat achieves 23.93% Pr@0.5 on RSREC task, noticeably lower than Qwen2.5-VL’s 28.66%. This result suggests that VLMs pretrained on large-scale generic datasets can exhibit strong cross-domain generalization even without access to remote sensing data. For RSRES task, we utilize outputs of VLMs as box prompts for SAM to get pixel-level masks. As we can see, Qwen2.5-VL still outperforms GeoChat by 0.8% in mIoU. These findings highlight the potential of generic VLMs as a solid foundation for zero-shot open-vocabulary RSVG, even without any domain-specific training.
@@
Refine Stage (Optional). Following prior work (liu2024rotated; dong2024cross), we additionally introduce an optional refinement stage to further improve the segmentation mask $\mathbf{M}$. We evaluate the effectiveness of several refinement methods, including DenseCRF and various prompt types for SAM. Our results show that box prompts yield the best performance, whereas point prompts are less effective.
@@
Datasets. We evaluate our method on two benchmarks tailored for RSVG: RRSIS-D (liu2024rotated) and RISBench (dong2024cross). Each sample comprises an RGB image, a referring expression, and a segmentation mask. Both datasets are extended from the RSREC dataset, inheriting its bounding box annotations, which allows for comprehensive evaluation of both REREC and RSRES tasks.
Implementation Details. Our RSVG-ZeroOV offers a zero-shot solution, requiring only an inference process, without the need for any training images or annotations. We employ the pre-trained Qwen2.5-VL (Qwen2.5-VL) and Stable Diffusion V1.4 (rombach2022high) as our VLM and DM, respectively. For Stable Diffusion, we use a guidance scale of 7.5 with 1,000 total diffusion steps, and perform DDIM-based sampling with 20 steps. The DDIM noise schedule follows a scaled linear progression from 0.00085 to 0.012. Following prior works (li2024language; liu2024rotated), we report mIoU, oIoU and Precision@X with $X\in\{0.3,0.5,0.7\}$. We set KK=7 for seed selection, $\tau$=0.3 for response thresholding, and $\alpha$=0.4 for binarization. All experiments are conducted on a single RTX4090 GPU.
@@
Comparison Results on RISBench. We further evaluate RSVG-ZeroOV on the more challenging RISBench, which features longer and more semantically complex referring expressions compared to RRSIS-D. As shown in Tab. 6, RSVG-ZeroOV (w/ Refine) demonstrates exceptional performance, establishing new SOTA benchmarks for Pr@0.5 and mIoU at 38.90%/38.87% and 31.03%/31.84% on both tasks, respectively. Notably, the proposed method outperforms the two zero-shot methods, OV-VG and GroundVLP by +18.8%/18.13% and +17.15%/16.27% in oIoU, respectively, indicating its superior ability to capture object structures and preserve overall shape integrity.
@@
Effect of self-attention resolutions. As shown in Tab. 8, we investigate the impact of varying resolutions of self-attention maps. We find that using multiple resolutions [32, 64] leads to the best performance, suggesting that multi-scale self-attention maps not only preserve high-resolution details but also effectively capture contextual semantics.
@@ Conclusion
In this paper, we introduce RSVG-ZeroOV, a zero-shot framework that integrates a VLM and a DM via attention maps to tackle the open-vocabulary remote sensing visual grounding. Without additional training, we design a novel overview-focus-evolve strategy to progressively enhance attention toward referred objects. Extensive experiments on existing benchmarks demonstrate that RSVG-ZeroOV achieves SOTA performance, outperforming existing weakly-supervised and zero-shot methods. To the best of our knowledge, RSVG-ZeroOV is the first zero-shot framework for remote sensing visual grounding. We hope this work can potentially inspire researchers to address open-vocabulary remote sensing tasks from a fresh perspective.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0117103, in part by the National Natural Science Foundation of China under Grants 62577041, in part by the Outstanding Youth Science Foundation of Shaanxi Province under Grant 2025JC-JCQN-083, in part by the Key Research and Development Program of Shaanxi Province under Grants 2025CY-YBXM-047 and 2024GX-YBXM-140.